# Impartial Games: a Challenge for Reinforcement Learning

### v5

@@ 3 Revisiting AlphaZero and LCZero
AlphaZero, while groundbreaking, was not released open-source. However, several community-driven self-play RL projects, such as LCZero for chess, Leela Zero and KataGo [42] for Go, and AobaZero  [^2] http://www.yss-aya.com/aobazero/index_e.html for shogi, have successfully replicated and extended the core principles of AlphaZero [10]. These projects leverage distributed computing, relying on contributions from a global community of enthusiasts, enabling the training of remarkably powerful game-playing agents. Notable open-source variants also include the Elf Go project [38] and a publicly released variant of AlphaZero [42].
AlphaZero and LCZero often approach position evaluation and move generation in ways that fundamentally differ from human grandmasters or traditional, hand-coded chess engines. Their revolutionary evaluations and surprising moves have profoundly impacted the chess community [28]. This section focuses on a highly trained version of LCZero333version: Last T60: 611246 (384 $\times$ 30), which utilizes a neural network architecture with 30 residual blocks and 384 filters, surpassing AlphaZero’s initial 20×25620$\times$ 25620 × 256 configuration. LCZero additionally incorporates Squeeze-and-Excitation layers into its residual blocks and supports endgame tablebases [21], potentially enabling it to exceed the original AlphaZero’s capabilities. Earlier LCZero versions (e.g., 20×25620$\times$ 25620 × 256 architecture) are generally weaker than these more advanced iterations.
For comparative analysis, we also utilize open-source Stockfish 14, an evolution of the traditional Stockfish engine (initially developed by Tord Romstad, Marco Costalba, and Joona Kiiski). Unlike its predecessor, Stockfish 14 integrates an Efficiently Updatable Neural Network (NNUE) for position evaluation [23], blending traditional search with neural network capabilities. While game-playing strength is commonly measured by Elo rating (detailed in Section 5.2.1), alternative metrics assessing move quality rather than just game outcomes have also been proposed [27]. As of August 7, 2021, Stockfish 14 held an Elo rating of 3555, Stockfish 8 3375, and LCZero 3336 on popular engine rating lists. LCZero, when run on comparable hardware to AlphaZero against Stockfish 8 with a fixed time control (1 minute per move), replicated AlphaZero’s impressive performance.
@@
As shown in Figure 3(d), with an increasing number of MCTS simulations, AlphaZero’s win probability initially showed a slight dip as it explored various defensive lines, before climbing again. Ultimately, after examining more than four million nodes, both engines yielded nearly identical evaluations, with win probabilities of 73.5% and 73.1% respectively. This demonstrates that LCZero effectively replicates AlphaZero’s capabilities in chess [28]. Even moves selected by LCZero without extensive search (i.e., greedily based on PV-NN evaluations) perform remarkably well, occasionally defeating strong human players.
@@
Despite being equipped with highly sophisticated PV-NNs, LCZero generally struggles to accurately evaluate positions with strong parity-related properties, including scenarios involving "zugzwang," "waiting moves," and "triangle maneuvers," which are common themes in chess endgames. In standard gameplay, these issues are often mitigated because the extensive MCTS compensates for initial PV-NN inaccuracies.
@@ 4 Two Levels of Mastery
In the domain of combinatorial game theory, a two-player game is considered solved if its optimal outcome can be predicted from any position, assuming both players make optimal moves [2]. Game solutions are typically categorized into three levels. An ultra-weak solution determines only whether the first player has a winning advantage from the initial position [39]. Given that self-play reinforcement learning (RL) agents need to generalize and make effective moves from arbitrary positions encountered during gameplay, the ultra-weak solution concept is less relevant to our investigation. Instead, we focus on the more comprehensive definitions of game solutions, as summarized in Table 1.
@@
Building upon these established definitions, we propose two distinct measures to assess the extent of mastery achieved by an RL algorithm: champion and expert. These measures characterize the depth and breadth of an agent’s understanding and strategic capability within a game, as delineated in Table 2.
@@
This distinction is of paramount importance for evaluating the true generalization capabilities of RL agents. The question posed by [18], regarding whether an RL agent truly learns "general knowledge applicable to any legal game state," directly pertains to its capacity to become an expert. A champion agent might possess a winning strategy from the initial state by effectively guiding the game into its "comfort zone," where its knowledge is deep and reliable. In contrast, an expert agent must consistently make optimal decisions across the entire relevant game graph, irrespective of whether the position is common or rare. While it is beyond the scope of this paper to provide a formal proof, it is intuitively plausible that in some games, becoming a champion without simultaneously becoming an expert is impossible, especially when the winning player cannot sufficiently constrain the opponent’s options to force the game into well-understood territories. A more rigorous theoretical analysis of the conditions under which champion-level play necessitates expert-level knowledge is an intriguing avenue for future research.
The concept of "problem chess," involving artificially composed chess positions designed to test specific tactical or strategic motifs, offers a relevant analogy. LCZero, like AlphaZero, is not explicitly trained to solve such artificial problems, and it is therefore unsurprising that it can struggle with them [21]. The primary objective of AlphaZero-style algorithms in chess, shogi, and Go has been to learn to play and win games from standard starting positions—a "champion" notion of mastery. Their success is measured by their ability to achieve superior game outcomes rather than by their capacity to find optimal moves in every conceivable (and potentially rare or adversarial) situation, which aligns with our "expert" definition. Our work explores the extent to which self-play RL, trained for champion-level performance, can truly generalize to expert-level play, particularly in games with a strong parity structure like Nim.
@@
Example 2: The [2,1,,1][2,1$\ldots$,1][ 2 , 1 , … , 1 ] Nim Board. Let $N\in\mathbb{N}$ and consider a Nim board configured as [2,1,1,,1][2,1,1$\ldots$,1][ 2 , 1 , 1 , … , 1 ], where there is one heap of 2 counters and $N$ heaps of 1 counter, with $N$ being a large number (e.g., $N=100$). An RL agent is likely to become a champion in this specific game configuration after relatively little training. It can quickly learn the optimal opening moves: specifically, to remove either 1 or 2 counters from the initial heap of 2, thus bringing the nim-sum to zero and achieving a P-position. Through self-play, the agent effectively memorizes these few crucial initial moves.
However, the task of selecting the best move in a general position derived from this board, represented as $[v_{1},v_{2},\ldots,v_{N+1}]$ where $v_{1}\in\{0,1,2\}$ and $v_{j}\in\{0,1\}$ for $j\in\{2,\ldots,N+1\}$, is fundamentally equivalent to evaluating the parity of the sub-configuration $[v_{2},\ldots,v_{N+1}]$. This involves computing the XOR sum of $N$ binary values. Due to the inherent challenges of neural networks in modeling parity functions, especially without a strong incentive to learn such a generalizable representation during self-play, a champion agent for this specific Nim board is not expected to simultaneously become an expert agent across all possible derivative states. This example underscores how an agent can achieve winning performance in specific initial conditions without developing a general, expert-level understanding of the underlying mathematical properties that govern all possible game states.
@@ 5 Reinforcement Learning for Nim
The AlphaZero-style learning paradigm has achieved unprecedented success in complex board games, with various open-source implementations available  [^4] https://github.com/suragnair/alpha-zero-general [^5] https://github.com/junxiaosong/AlphaZero_Gomoku and specialized RL libraries offering optimized frameworks [19, 20]. However, for our specific diagnostic analysis of AlphaZero-style algorithms on impartial games, these existing solutions lacked critical functionalities. Our objectives necessitated granular control over the training process and detailed insights into the network’s internal representations, including, but not limited to, the real-time calculation of agent Elo ratings against its ancestors and, crucially, the precise evaluation of the policy network’s output accuracy against the known optimal moves derived from the nim-sum function as training progresses. These diagnostic capabilities are pivotal for understanding the learning dynamics in domains where optimal play is mathematically defined.
This necessitated the implementation of a custom AlphaZero-style self-play algorithm. Our implementation leverages PyTorch [25] for neural network training and Ray [22] for efficient parallelization of MCTS simulations. We made considerable efforts to ensure that our implementation rigorously replicates the architectural and algorithmic details of the original AlphaZero algorithms as described in [35] and [36]. While maintaining this high fidelity, necessary adaptations to the neural network architectures were made to accommodate the specific input and output representations of Nim, which will be elaborated in detail in Section 5.1.
To contextualize this, consider our analysis of a specific Chess position (Fig. 5) which subtly mirrors a rudimentary parity (nim-like) configuration. Our investigation revealed that even the highly optimized and extensively trained LCZero, which for all practical purposes embodies the AlphaZero paradigm, struggled to accurately assess this particular position. This observation provides a compelling precursor to our central argument: in Nim, where the winning strategy is directly and inextricably linked to the parity function (nim-sum), this issue becomes dramatically more pronounced. The core difficulty lies in the generic neural network’s struggle to implicitly learn a perfect parity function from noisy, self-play generated data, especially as the input dimensionality increases. This inherent representational bottleneck for combinatorial logic, distinct from the complexity of typical board game patterns, is what we aim to thoroughly explore in the subsequent sections.
@@ 5.1 Implementing an AlphaZero-style algorithm for Nim
The AlphaZero algorithm initiates training from a blank slate state, with its sole prior knowledge being the game’s fundamental rules [36]. Its architecture relies on three interdependent components: the policy network, the value network, and the MCTS. The policy network, typically a deep neural network, outputs a probability distribution $P(s,a)$ over all legal actions $a\in\mathcal{A}$ for a given state $s$. Probabilities for illegal actions are set to zero, and the remaining probabilities are re-normalized. This network serves to narrow the search space by prioritizing actions with a high likelihood of leading to favorable outcomes, thereby reducing the MCTS’s breadth. Concurrently, the value network outputs a scalar value $v\in[-1,1]$, providing an estimate of the expected outcome from state $s$. A higher value indicates a greater probability of the current player winning, and vice versa. By predicting the value of leaf nodes, the value network effectively prunes the search tree’s depth, as a full rollout to the game’s conclusion is computationally expensive and its true outcome (1 for win, -1 for loss) is only definitively known at termination.
@@
The training targets for the value network are derived from the actual outcomes of self-play games. If the agent wins a game, all encountered positions within that game are labeled as winning states (target value +1), and conversely, losing positions are labeled as -1. This self-generated labeling process is susceptible to noise, especially during the early stages of training when the PV-NN is untrained or weakly trained. Such noisy labels pose a significant challenge for neural networks, particularly in learning abstract functions like parity on long bitstrings [43]. Consequently, this inherent difficulty extends directly to the PV-NN’s ability to accurately evaluate positions in Nim, where the optimal strategy is fundamentally linked to the nim-sum.
@@
Recognizing that Long Short-Term Memory (LSTM) networks are theoretically capable of modeling perfect parity functions due to their ability to learn long-range dependencies in sequential data [11], we employed an LSTM-based architecture for our PV-NN. The policy and value networks share a common LSTM layer, which then branches into two separate heads: a policy head and a value head. For Nim games with 5, 6, and 7 heaps, the shared LSTM layer used a hidden size of 128. While it is intuitive to scale network depth with increasing board size, we empirically found that employing a larger number of LSTM layers was detrimental to performance, often destabilizing the training process. The output of the shared LSTM layer feeds into the two heads. The policy head consists of 25, 36, and 49 nodes for 5, 6, and 7 heaps Nim respectively (matching the action space), with its logits passed through a softmax function to yield a probability distribution over actions. The value head comprises a single node, outputting a scalar value passed through a tanh activation function to ensure its range is confined to [1,1][-1,1][ - 1 , 1 ].
@@
The constant $c_{put}$ controls the balance between exploitation and exploration. While AlphaGo [34], AlphaGo Zero [35], and AlphaZero [36] left $c_{put}$ unspecified or tuned it empirically for specific games, we found its value significantly impacted performance on Nim. Setting $c_{put}$ too low excessively discourages exploration, while setting it too high diminishes the action value’s influence, impairing search depth effectiveness. Initial attempts with values common in other domains, such as $c_{put}=\{1,1.5,2,3\}$ (where $c_{put}=1.5$ yielded satisfactory results for Elf OpenGo [38]), performed poorly for Nim. Consequently, we adopted a dynamically adjusted exploration term inspired by MuZero [30]:
@@
[…]
@@
After completing a predefined number of simulations for each move, the MCTS returns a robust probability distribution $\bm{\pi}(\textbf{a}|s)$ over all actions, from which an action is sampled for execution. This distribution is calculated as:
@@
Here, $\tau$ represents the temperature parameter, which modulates the exploration/exploitation balance. We set $\tau=1$ for the first 3 moves of a game, allowing action selection to be proportional to their visit counts, thus encouraging broader exploration in the opening. For subsequent moves, $\tau=0$ is used, ensuring that the action with the highest visit count is deterministically chosen (pure exploitation). During the evaluation phase, $\tau$ is consistently set to 0 for all moves. These posterior probabilities $\bm{\pi}(\textbf{a}|s)$ (derived from MCTS simulations guided by the policy network’s priors) serve as the training targets for the policy network, aiming to refine its initial predictions.

### v6

@@ 3 Revisiting AlphaZero and LCZero
AlphaZero, while groundbreaking, was not released as open-source. However, several community-driven self-play RL projects, such as LCZero for chess, Leela Zero and KataGo (Wu, 2019) for Go, and AobaZero [^2] AobaZero: http://www.yss-aya.com/aobazero/index_e.html for shogi, have successfully replicated and extended the core principles of AlphaZero (Cazenave et al., 2020). These projects leverage distributed computing, relying on contributions from a global community of enthusiasts, enabling the training of remarkably powerful game-playing agents. Notable open-source variants also include the Elf Go project (Tian et al., 2019) and a publicly released variant of AlphaZero (Wu, 2019).
AlphaZero and LCZero often approach position evaluation and move generation in ways that fundamentally differ from human grandmasters or traditional, hand-coded chess engines. Their revolutionary evaluations and surprising moves have profoundly impacted the chess community (Sadler and Regan, 2019). This section focuses on a highly trained version of LCZero [^3] Version: Last T60: 611246 (384 $\times$ 30), which utilises a neural network architecture with 30 residual blocks and 384 filters, surpassing AlphaZero’s initial $20\times 256$ configuration. LCZero additionally incorporates Squeeze-and-Excitation layers into its residual blocks and supports endgame tablebases (Maharaj et al., 2021), potentially enabling it to exceed the original AlphaZero’s capabilities. Earlier LCZero versions (e.g., $20\times 256$ architecture) are generally weaker than these more advanced iterations.
For comparative analysis, we also utilise open-source Stockfish 14, an evolution of the traditional Stockfish engine (initially developed by Tord Romstad, Marco Costalba, and Joona Kiiski). Unlike its predecessor, Stockfish 14 integrates an Efficiently Updatable Neural Network (NNUE) for position evaluation (Nasu, 2018), blending traditional search with neural network capabilities. While game-playing strength is commonly measured by Elo rating (detailed in Section 5.2.1), alternative metrics assessing move quality rather than just game outcomes have also been proposed (Regan and Haworth, 2011). As of August 7, 2021, Stockfish 14 held an Elo rating of 3555, Stockfish 8 3375, and LCZero 3336 on popular engine rating lists. LCZero, when run on comparable hardware to AlphaZero against Stockfish 8 with a fixed time control (1 minute per move), replicated AlphaZero’s impressive performance.
@@
As shown in Figure 3(d), with an increasing number of MCTS simulations, AlphaZero’s win probability initially showed a slight dip as it explored various defensive lines, before climbing again. Ultimately, after examining more than four million nodes, both engines yielded nearly identical evaluations, with win probabilities of 73.5% and 73.1%, respectively. This demonstrates that LCZero effectively replicates AlphaZero’s capabilities in chess (Sadler and Regan, 2019). Even moves selected by LCZero without extensive search (i.e., greedily based on PV-NN evaluations) perform remarkably well, occasionally defeating strong human players.
@@
Despite being equipped with highly sophisticated PV-NNs, LCZero generally struggles to accurately evaluate positions with strong parity-related properties, including scenarios involving "zugzwang," "waiting moves," and "triangle manoeuvres," which are common themes in chess endgames. In standard gameplay, these issues are often mitigated because the extensive MCTS compensates for initial PV-NN inaccuracies.
@@ 4 Two levels of mastery
In the domain of combinatorial game theory, a two-player game is considered solved if its optimal outcome can be predicted from any position, assuming both players make optimal moves (Allis and others, 1994). Game solutions are typically categorised into three levels. An ultra-weak solution determines only whether the first player has a winning advantage from the initial position (Van Den Herik et al., 2002). Given that self-play RL agents need to generalise and make effective moves from arbitrary positions encountered during gameplay, the ultra-weak solution concept is less relevant to our investigation. Instead, we focus on the more comprehensive definitions of game solutions, as summarised in Table 1.
@@
Building upon these established definitions, we propose two distinct measures to assess the extent of mastery achieved by an RL algorithm: champion and expert. These measures characterise the depth and breadth of an agent’s understanding and strategic capability within a game, as delineated in Table 2.
@@
This distinction is of paramount importance for evaluating the true generalisation capabilities of RL agents. The question posed by Lan et al. (2022), regarding whether an RL agent truly learns "general knowledge applicable to any legal game state," directly pertains to its capacity to become an expert. A champion agent might possess a winning strategy from the initial state by effectively guiding the game into its "comfort zone," where its knowledge is deep and reliable. In contrast, an expert agent must consistently make optimal decisions across the entire relevant game graph, irrespective of whether the position is common or rare. While it is beyond the scope of this paper to provide a formal proof, it is intuitively plausible that in some games, becoming a champion without simultaneously becoming an expert is impossible, especially when the winning player cannot sufficiently constrain the opponent’s options to force the game into well-understood territories. A more rigorous theoretical analysis of the conditions under which champion-level play necessitates expert-level knowledge is an intriguing avenue for future research.
The concept of "problem chess," involving artificially composed chess positions designed to test specific tactical or strategic motifs, offers a relevant analogy. LCZero, like AlphaZero, is not explicitly trained to solve such artificial problems, and it is therefore unsurprising that it can struggle with them (Maharaj et al., 2021). The primary objective of AlphaZero-style algorithms in chess, shogi, and Go has been to learn to play and win games from standard starting positions—a "champion" notion of mastery. Their success is measured by their ability to achieve superior game outcomes rather than by their capacity to find optimal moves in every conceivable (and potentially rare or adversarial) situation, which aligns with our "expert" definition. Our work explores the extent to which self-play RL, trained for champion-level performance, can truly generalise to expert-level play, particularly in games with a strong parity structure like Nim.
@@
Example 2: The $2,1,\ldots,1]$ Nim Board. Let $N\in\mathbb{N}$ and consider a Nim board configured as $2,1,1,\ldots,1]$, where there is one heap of 2 counters and $N$ heaps of 1 counter, with $N$ being a large number (e.g., $N=100$). An RL agent is likely to become a champion in this specific game configuration after relatively little training. It can quickly learn the optimal opening moves: specifically, to remove either 1 or 2 counters from the initial heap of 2, thus bringing the nim-sum to zero and achieving a P-position. Through self-play, the agent effectively memorises these few crucial initial moves.
However, the task of selecting the best move in a general position derived from this board, represented as $v_{1},v_{2},\ldots,v_{N+1}]$ where $v_{1}\in\{0,1,2\}$ and $v_{j}\in\{0,1\}$ for $j\in\{2,\ldots,N+1\}$, is fundamentally equivalent to evaluating the parity of the sub-configuration $v_{2},\ldots,v_{N+1}]$. This involves computing the XOR sum of $N$ binary values. Due to the inherent challenges of neural networks in modelling parity functions, especially without a strong incentive to learn such a generalizable representation during self-play, a champion agent for this specific Nim board is not expected to simultaneously become an expert agent across all possible derivative states. This example underscores how an agent can achieve winning performance in specific initial conditions without developing a general, expert-level understanding of the underlying mathematical properties that govern all possible game states.
@@ 5 Reinforcement learning for Nim
The AlphaZero-style learning paradigm has achieved unprecedented success in complex board games, with various open-source implementations available [^4] AlphaZero-general: https://github.com/suragnair/alpha-zero-general; AlphaZero_Gomoku: https://github.com/junxiaosong/AlphaZero_Gomoku and specialized RL libraries offering optimized frameworks (Lanctot et al., 2019; Liang et al., 2018). However, for our specific diagnostic analysis of AlphaZero-style algorithms on impartial games, these existing solutions lacked critical functionalities. Our objectives necessitated granular control over the training process and detailed insights into the network’s internal representations, including, but not limited to, the real-time calculation of agent Elo ratings against its ancestors and, crucially, the precise evaluation of the policy network’s output accuracy against the known optimal moves derived from the nim-sum function as training progresses. These diagnostic capabilities are pivotal for understanding the learning dynamics in domains where optimal play is mathematically defined.
This necessitated the implementation of a custom AlphaZero-style self-play algorithm. Our implementation leverages PyTorch (Paszke et al., 2019) for neural network training and Ray (Moritz et al., 2018) for efficient parallelisation of MCTS simulations. We made considerable efforts to ensure that our implementation rigorously replicates the architectural and algorithmic details of the original AlphaZero algorithms as described in Silver et al. (2017) and Silver et al. (2018). While maintaining this high fidelity, necessary adaptations to the neural network architectures were made to accommodate the specific input and output representations of Nim, which will be elaborated in detail in Section 5.1.
@@ 5.1 Implementing an AlphaZero-style algorithm for Nim
The AlphaZero algorithm initiates training from a blank slate state, with its sole prior knowledge being the game’s fundamental rules (Silver et al., 2018). Its architecture relies on three interdependent components: the policy network, the value network, and the MCTS. The policy network, typically a deep neural network, outputs a probability distribution $P(s,a)$ over all legal actions $a\in\mathcal{A}$ for a given state $s$. Probabilities for illegal actions are set to zero, and the remaining probabilities are re-normalised. This network serves to narrow the search space by prioritising actions with a high likelihood of leading to favourable outcomes, thereby reducing the MCTS’s breadth. Concurrently, the value network outputs a scalar value $v\in[-1,1]$, providing an estimate of the expected outcome from state $s$. A higher value indicates a greater probability of the current player winning, and vice versa. By predicting the value of leaf nodes, the value network effectively prunes the search tree’s depth, as a full rollout to the game’s conclusion is computationally expensive and its true outcome (1 for win, -1 for loss) is only definitively known at the end of the game.
@@
The training targets for the value network are derived from the actual outcomes of self-play games. If the agent wins a game, all encountered positions within that game are labelled as winning states (target value +1), and conversely, losing positions are labelled as -1. This self-generated labelling process is susceptible to noise, especially during the early stages of training when the PV-NN is untrained or weakly trained. Such noisy labels pose a significant challenge for neural networks, particularly in learning abstract functions like parity on long bitstrings (Zhou and Riis, 2023). Consequently, this inherent difficulty extends directly to the PV-NN’s ability to accurately evaluate positions in Nim, where the optimal strategy is fundamentally linked to the nim-sum.
@@
Recognising that Long Short-Term Memory (LSTM) networks are theoretically capable of modelling perfect parity functions due to their ability to learn long-range dependencies in sequential data (Cohen-Karlik et al., 2020), we employed an LSTM-based architecture for our PV-NN. The policy and value networks share a common LSTM layer, which then branches into two separate heads: a policy head and a value head. For Nim games with 5, 6, and 7 heaps, the shared LSTM layer used a hidden size of 128. While it is intuitive to scale network depth with increasing board size, we empirically found that employing a larger number of LSTM layers was detrimental to performance, often destabilising the training process. The output of the shared LSTM layer feeds into the two heads. The policy head consists of 25, 36, and 49 nodes for 5, 6, and 7 heaps Nim, respectively (matching the action space), with its logits passed through a softmax function to yield a probability distribution over actions. The value head comprises a single node, outputting a scalar value passed through a tanh activation function to ensure its range is confined to [1,1][-1,1].
@@
The constant $c_{put}$ controls the balance between exploitation and exploration. While AlphaGo (Silver et al., 2016), AlphaGo Zero (Silver et al., 2017), and AlphaZero (Silver et al., 2018) left $c_{put}$ unspecified or tuned it empirically for specific games, we found its value significantly impacted performance on Nim. Setting $c_{put}$ too low excessively discourages exploration, while setting it too high diminishes the action value’s influence, impairing search depth effectiveness. Initial attempts with values common in other domains, such as $c_{put}=\{1,1.5,2,3\}$ (where $c_{put}=1.5$ yielded satisfactory results for Elf OpenGo (Tian et al., 2019)), performed poorly for Nim. Consequently, we adopted a dynamically adjusted exploration term inspired by MuZero (Schrittwieser et al., 2020):
@@
Here, $\eta_{a}$ is sampled from a Dirichlet distribution Dir($\alpha$), with $\alpha$ set to 0.35. The constant $\epsilon$ is set to 0.25 during training but to 0 during evaluation to eliminate the noise effect. The typical $\alpha$ values used in AlphaZero for Chess, Shogi, and Go are 0.3, 0.15, and 0.003, respectively. Given that the average number of legal moves in the Nim variants we study is generally lower than in Chess, we opted for a higher $\alpha$ value of 0.35 to promote more diverse exploration relative to the smaller action space. While theoretically $\alpha=0.5$ might be optimal for maximising entropy, our empirical fine-tuning found $\alpha=0.35$ to yield better practical outcomes.
@@
After completing a predefined number of simulations for each move, the MCTS returns a probability distribution $\bm{\pi}(\textbf{a}|s)$ over all actions, from which an action is sampled for execution. This distribution is calculated as:
@@
Here, $\tau$ represents the temperature parameter, which modulates the exploration/exploitation balance. We set $\tau=1$ for the first 3 moves of a game, allowing action selection to be proportional to their visit counts, thus encouraging broader exploration in the opening. For subsequent moves, $\tau=0$ is used, ensuring that the action with the highest visit count is deterministically chosen (pure exploitation). During the evaluation phase, $\tau$ is consistently set to 0 for all moves. These posterior probabilities $\bm{\pi}(\textbf{a}|s)$ serve as the training targets for the policy network, aiming to refine its initial predictions.